# Unveiling Language-Specific Features in Large Language Models via Sparse Autoencoders

### v1

@@ Appendix D Additional Results for Directional Ablation
Additional results for directional ablation for 3 different LLMs are demonstrated in Figure 13-15. The results are similar to those in Figure 5, where Ablating language-specific features has a much larger impact on the CE loss of texts in target language compared to texts in other languages.
@@ Appendix E Additional Results for Code-Switching
Figure 25: The mean activation values for various feature with Spanish prefix for Gemma 2 2B.
Figure 26: The mean activation values for various feature with Japanese prefix for Gemma 2 2B.
Figure 27: The mean activation values for various feature with Thai prefix for Gemma 2 2B.
Figure 28: The mean activation values for various feature with Spanish prefix for Gemma 2 9B.
Figure 29: The mean activation values for various feature with Japanese prefix for Gemma 2 9B.
Figure 30: The mean activation values for various feature with Thai prefix for Gemma 2 9B.
Figure 31: The mean activation values for various feature with Spanish prefix for Llama-3.1-8B.
Figure 32: The mean activation values for various feature with Japanese prefix for Llama-3.1-8B.
Figure 33: The mean activation values for various feature with Thai prefix for Llama-3.1-8B.

### v2

@@ 8 Conlusion
Acknowledgements
This research was also supported by the advanced computing resources provided by the Supercomputing Center of the USTC.
@@ Appendix D Additional Results for Directional Ablation
Additional results for directional ablation for 3 different LLMs are demonstrated in Figure 13-15. The results are similar to those in Figure 5, where ablating language-specific features has a much larger impact on the CE loss of texts in the target language compared to texts in other languages.
@@ Appendix E Additional Results for Code-Switching
Figure 25: The mean activation values for various features with Spanish prefix for Gemma 2 2B.
Figure 26: The mean activation values for various features with Japanese prefix for Gemma 2 2B.
Figure 27: The mean activation values for various features with Thai prefix for Gemma 2 2B.
Figure 28: The mean activation values for various features with Spanish prefix for Gemma 2 9B.
Figure 29: The mean activation values for various features with Japanese prefix for Gemma 2 9B.
Figure 30: The mean activation values for various features with Thai prefix for Gemma 2 9B.
Figure 31: The mean activation values for various features with Spanish prefix for Llama-3.1-8B.
Figure 32: The mean activation values for various features with Japanese prefix for Llama-3.1-8B.
Figure 33: The mean activation values for various features with Thai prefix for Llama-3.1-8B.